### v1

@@ 1.7 An Overview of Necessities for Reliable AI
The conditions for reliable AI are still unclear and are addressed incongruously by many institutions and groups, despite contentious social debates over it. The principles of Fairness, Responsibility, Accuracy, Verifiability, and Accountability in Machine Learning (FAT-ML) include accountability, explainability, verifiability, and fairness at a global level [32]. Among the numerous needs under review, explainability, fairness, privacy, and robustness will all be covered in this study (Table 1).
@@
2 Trustworthy XAI Vs AI
A paradigm change in the field of artificial intelligence (AI) has been brought about with the introduction of Trustworthy Explainable AI (XAI). Oftentimes, conventional AI systems operate as opaque black boxes, making it challenging for users to comprehend the decision-making process. On the other hand, Trustworthy XAI tackles significant concerns surrounding the adoption of AI by focusing on accountability, interpretability, and transparency. Reputable XAI seeks to build consumers’ confidence by providing clear justifications for its choices. Users are able to evaluate the fairness and dependability of AI-driven results because to this transparency. While typical AI systems are capable of producing precise forecasts or suggestions, they do not have the openness required to establish credibility.
The way that reliable XAI and traditional AI make decisions is what sets them apart from one another. Although AI systems have the potential to produce precise forecasts or suggestions, Reputable XAI highlights the necessity of providing justifications that clarify the processes that lead to these results. Users may evaluate the fairness and dependability of AI-driven outcomes thanks to the clear justifications that trustworthy XAI systems give for their decisions.
Furthermore, trustworthy XAI takes into account more ethical factors than just explanations. Incorporating the Fairness, Accountability, and Transparency (FAT) principles into AI development procedures, guarantees that AI systems abide by moral and legal requirements. Trustworthy XAI seeks to reduce biases, discrimination, and other possible negative effects of AI technology by placing a high priority on ethical norms. An approach to AI development known as trustworthy AI places a high value on user safety and openness. Since no model is flawless, trustworthy AI developers take care to explain to clients and the wider public how the technology was developed, its intended applications, and its limitations.
@@ 2 Trustworthy XAI Vs AI
Transparency Systems using artificial intelligence ought to be transparent and traceable The right to get clarification Recital 71
@@ 3 Applications of Trustworthy XAI
Authentic Explainable Artificial Intelligence (XAI) has numerous uses in sectors where accountability, interpretability, and transparency are essential. XAI can provide an explanation for a diagnosis or therapy recommendation in medical diagnosis and recommendation systems. Financial institutions can employ XAI for risk assessment, fraud detection, and credit scoring. XAI can help attorneys with contract analysis, lawsuit prediction, and legal research. In autonomous vehicles, XAI plays a significant role in providing context for the decisions made by the AI systems, particularly in high-stakes scenarios such as accidents or unanticipated roadside incidents. XAI can be applied to process optimization, predictive maintenance, and quality control in manufacturing settings. By offering justifications for automated responses or suggestions in chatbots and virtual assistants, XAI can improve customer service. By providing an explanation for the recommendations and assessments made by adaptive learning systems, XAI can help with individualized learning. By providing an explanation for the recommendations and assessments made by adaptive learning systems, XAI can help with individualized learning. We shall concentrate on a few particular applications in this section and go into detail about them.
3.1 Application of Trustworthy XAI in Medical Science
The field of artificial intelligence (AI) is rapidly growing on a global scale. The potential uses of artificial intelligence in healthcare are a hot topic for research [33]. There are many opportunities to use AI technology in the healthcare sector, where people’s lives and well-being are in danger, because of its essential relevance and the enormous quantity of digital medical data that have been gathered [34]. Artificial intelligence (AI) has made it possible to accomplish tasks quickly that were before unfeasible for traditional technologies. Trustworthy AI is a huge concern these days. Since incidents involving AI-powered chatbots such as Tay and Iruda [35], there has been an increase in interest in the topic of whether an AI’s judgment and decision-making system is reliable. The credibility of AI in the medical and healthcare areas requires more investigation. Clinical decision support systems (CDSS) in the medical field use AI technology to aid with important medical duties such as diagnosis and therapy planning [36]. Misuse can have serious consequences in areas where lives are at stake, even if the scope of use is limited to assisting healthcare practitioners. False alarms, for example, which occur often in scenarios involving urgent patients, may exhaust medical personnel.
The study [37] adds significantly to the field of medical skin lesion diagnostics in a number of ways. Before anything else, it adapts an existing eXplainable Artificial Intelligence (XAI) technique to increase user confidence and trust in AI decision-making systems. This modification entails describing an AI model that is skilled at differentiating between different kinds of skin lesions. Synthetic exemplar and counter-exemplar images are used to create explanations that illustrate the important characteristics that influence classification choices. This research [37] is based on training a deep learning classifier with the ISIC 2019 dataset using the ResNet architecture. This enables practitioners to use the explanations offered to reason effectively. All things considered, the study’s original contributions are found in the way it refined and assessed the XAI technique in an actual medical setting, examined the latent space and carried out an extensive user study to gauge the efficacy of the explanations, especially among subject matter experts.
This research paper [38] addresses the challenge of recognizing brain tumors in MRI images by merging two powerful algorithms: fuzzy C-means (the FCM method) and Artificial Neural Network (ANN). The authors want to increase the segmentation process’s interpretability as well as the accuracy of tumor identification by merging these techniques. Their main objective is to improve medical decision-support system tools so that physicians can diagnose brain tumors more accurately. This discovery has two main benefits: first, it improves the ability to identify brain cancers in medical imaging with more accuracy, which is important for early diagnosis and treatment. Second, the researchers make the decisions made by their models more transparent and intelligible to patients and medical experts by integrating explainable AI principles into the segmentation process. In the end, this improved interpretability raises the general level of reliability and acceptance of AI-driven systems for medical picture segmentation in clinical settings.
In another research [39], the field of computational pathology—which uses machine learning and artificial intelligence (AI) to diagnose whole slide images (WSIs)—is discussed. Because artificial intelligence is opaque, there are doubts regarding its reliability despite the technology’s potential to improve efficiency and accuracy. The paper suggests employing explainable AI (xAI) techniques, which can shed light on the choices made by AI algorithms, to allay these worries. Computational pathology systems become more clear and dependable with the addition of xAI, especially when it comes to crucial activities like pathology diagnosis. Additionally, it presents HistoMapr-Breast, a software program with xAI capabilities intended for breast core biopsies.
In a recent research [40], it is discussed how crucial it is to guarantee the correctness and resilience of AI-based systems in the healthcare industry, especially with regard to their interpretability and defense against hostile attacks. As AI systems are used in medical contexts more frequently, it is important to confirm that the predictions they are generating are based on accurate features. Numerous model interpretability and explainability techniques have been put out in an effort to overcome this. This work shows that even with strong training, adversarial attacks can affect the explainability of a model. The authors also introduce two attack classifiers: one to determine the type of attack and another to differentiate between benign and opposed inputs.
This research paper [41] studies the rising area of explainable machine learning in cardiology. It addresses challenges with the interpretability of complicated prediction models and their ramifications for key healthcare choices. This study delves into the underlying principles and tactics of explainable machine learning, providing cardiologists with a greater grasp of the approach’s strengths and limitations. The research intends to help decision-making processes in clinical settings by providing a rule of thumb for deciding between interpretable and black-box models. This will ultimately improve patient outcomes while maintaining accountability and transparency in model predictions. Figure 9 was created by training a single new decision tree based on the random forest model predictions. The global tree diagram depicts the overall operation of the random forest. Individual predictions can then be investigated by tracking the patient’s progress through the global tree. This type of explanation has the advantage of making it easy to understand both the overall operation of the model and the reasoning behind each particular forecast. Decision trees are ideal for disciplines such as cardiology because they provide rule-based reasoning similar to clinical decision criteria.
Figure 9: A random forest-based model for heart disease prediction is described, using both local and global decision trees. Because decision trees are evaluated from top to bottom, the global diagram shows that the model begins by deciding if patients’ thallium stress test results are normal. The model investigates the patient’s ST depression if the thallium stress test discloses an issue, and so on. The local graphic depicts the path a patient took to fall from the tree, showing the reasoning for each individual prognosis. The patient was less than 54.5 years old, had a maximum heart rate of higher than, and had a normal thallium stress test result, indicating a very low chance of cardiac disease [41].
@@ 3.1 Application of Trustworthy XAI in Medical Science
Figure 10 depicts the LIME explanations for our heart failure model’s two local predictions. The authors demonstrate how these predictions are integrated as a clinical decision support tool in Epic, an electronic health record created with doctors in mind (Epic Systems Corporation, Verona, Wisconsin, USA).37 This type of explainability approach has the advantage of providing detailed explanations for the clinical factors that impact each prediction. It is noteworthy that this kind of explanation can be incorporated into an EHR, as this could enhance a black-box model’s actionability by allowing forecasts and clear explanations to be effortlessly integrated into clinical workflow.
@@
Reliable medical AI implementation and widespread usage need substantial research and social consensus on prerequisites such as explainability, fairness, privacy protection, and robustness [32]. Optimized requirements and standards must be satisfied in any therapeutic environment where AI is used, and these requirements and standards need to be updated often. Furthermore, laws must be put in place specifying who is in charge in the event of a medical AI-related mishap or accident—designers, researchers, medical personnel, or patients [42].
@@ 3.2 Explainability and Interpretability of Autonomous Systems
Explainability and interpretability in the context of autonomous systems refer to the ability to understand and make sense of the systems’ decisions and behaviors. Explainability refers to an autonomous system’s ability to provide clear arguments for its decisions and behaviors [43]. It is critical for increasing acceptance and confidence in AI systems, particularly in areas such as banking, healthcare, and autonomous autos. While explainability and interpretability are closely connected, interpretability places more emphasis on the capacity to comprehend the internal workings and procedures of the autonomous system [44]. An interpretable system provides users with insight into the elements and criteria considered while making decisions, allowing them to comprehend how the system came to its findings.
@@
The research paper The research article [18] focuses on trust and dependability in autonomous systems. Autonomous systems have the potential for system operation, rapid information dissemination, massive data processing, working in hazardous environments, operating with greater resilience and tenacity than humans, and even astronomical examination [45], [46]. Following years of research and development, today’s automated technologies represent the peak of progress in computer recognition, responsive systems, user-friendly interface design, and sensing automation. According to [43], the global market for automotive intelligent hardware, operations, and innovation is expected to rise from \$1.25 billion in 2017 to \$28.5 billion by 2025. According to Intel’s research of the predicted advantages of autonomous cars, employing these advances on public roads would decrease commute time by 250 million hours annually and save over 500,000 lives in the United States alone between 2035 and 2045 [43]. Modern automobiles employ artificial intelligence (AI) for several tasks, including intelligent cruise control, automatic driving and parking, and blind-spot identification (Figure 111).
@@
Authors [18] describe the challenges of autonomous systems, like, people sometimes tend to be overly excited about the potential of new ideas and ignore, or at least appear to be unaware of, the potential drawbacks of cutting-edge developments. Even in the early stages of robotics and autonomous system implementation, humanity preferred to put up with faulty goods and services, but they have gradually come to understand the importance of trustworthy and dependable autonomous systems [zablocki2022explainability]. Numerous examples have demonstrated how operators’ use of automation is greatly impacted by trustworthiness.
@@
Figure 11: An automated vehicle that provides a valid and understandable rationale for the decision it made at that particular instant, acting as the archetypal example of XAI in automated driving [18].
@@
When artificial intelligence (AI) has become prevalent in autonomous vehicle (AV) operations, user trust has been identified as a major issue that is essential to the success of these operations. Explainable artificial intelligence (XAI), which calls for the AI system to give the user explanations for every decision it makes, is a viable approach to fostering user trust for such integrated AI-based driving systems [47]. This work develops explainable Deep Learning (DL) models to improve trustworthiness in autonomous driving systems, driven by the need to improve user trust and the potential of innovative XAI technology in addressing such requirements. The main concept of this [47] research is to frame the decision-making process of autonomous vehicles (AVs) as an image-captioning task, generating textual descriptions of driving scenarios to serve as understandable explanations for humans. The proposed multi-modal deep learning architecture (Shown in Figure 12), based on Transformers, effectively models the correlation between images and language, generating meaningful descriptions and driving actions. Its contributions lie in formulating the traditional AV decision process for explainability, developing a fully Transformer-based model for generating descriptions and actions, and demonstrating superior performance over baseline models. The outcome is a model that enhances user trust, provides insights for AV developers, and offers superior interpretability through its attention mechanisms and end-to-end goal induction.
@@
This research [48] aims to investigate the integration of explainable artificial intelligence (XAI) into autonomous vehicular systems to improve transparency and human trust. It delves into the functioning of multiple inner vehicle modules, emphasizing the importance of understanding the vehicle’s decision-making processes for user credibility and reliability. The main contribution lies in introducing XAI to the domain of autonomous vehicles, showcasing its role in fostering trust, and highlighting advancements through comparative analysis. The output comprises the creation of visual explanatory techniques and an intrusion detection classifier, which show considerable advances over previous work in terms of transparency and safety in autonomous transportation systems.

### v2

@@ 1.8 An Overview of Necessities for Reliable AI
Despite heated societal discussions, the requirements for trustworthy AI remain ambiguous and are handled inconsistently by numerous organizations and organizations. Globally, accountability, explainability, verifiability, and fairness are all part of the Fairness, Responsibility, Accuracy, Verifiability, and Accountability in Machine Learning (FAT-ML) principles [33]. Explainability, fairness, privacy, and robustness are just a few of the many needs that will be examined in this study (Table 1).
@@
2 XAI Vs AI
XAI improves AI systems by focusing on transparency, clear explanations, and accountability. It offers understandable reasons for decisions, which helps users trust the system and makes it easier to assess fairness compared to traditional ”black box” methods.
The main difference between reliable XAI and traditional AI is how they make decisions. While AI can give accurate forecasts or suggestions, reliable XAI emphasizes the need to explain the steps that lead to these results. Clear explanations from XAI systems allow users to judge the fairness and reliability of AI-generated outcomes.
To improve security and maintain transparency in AI-driven cybersecurity, we need to integrate XAI into Zero Trust Architecture (ZTA). When explainability methods clarify why certain decisions are made, people can better understand and trust the AI-driven access control and behavioral analytics in ZTA. As we face compliance and operational challenges, future cybersecurity frameworks will rely more on AI automation. It will be essential to ensure that these AI systems can be easily explained [34].
XAI focuses on more than just providing explanations; it also considers ethical issues. AI development processes that follow the principles of Fairness, Accountability, and Transparency (FAT) help ensure that AI systems meet ethical and legal standards. By prioritizing ethical standards, XAI aims to reduce biases, discrimination, and other harmful effects of AI technology. Trustworthy AI is an approach that emphasizes user safety and transparency. Responsible AI developers clearly explain to clients and the public how the technology works, what it is meant for, and its limitations, since no model is perfect.
@@ 2 XAI Vs AI
Transparency Systems using AI ought to be transparent and traceable The right to get clarification Recital 71
@@ 3 Applications of XAI
Authentic XAI has numerous uses in sectors where accountability, interpretability, and transparency are essential. XAI can provide an explanation for a diagnosis or therapy recommendation in medical diagnosis and recommendation systems. Financial institutions can employ XAI for risk assessment, fraud detection, and credit scoring. XAI can help attorneys with contract analysis, lawsuit prediction, and legal research. In autonomous vehicles, XAI plays a significant role in providing context for the decisions made by the AI systems, particularly in high-stakes scenarios such as accidents or unanticipated roadside incidents. XAI can be applied to process optimization, predictive maintenance, and quality control in manufacturing settings. By offering justifications for automated responses or suggestions in chatbots and virtual assistants, XAI can improve customer service. By providing an explanation for the recommendations and assessments made by adaptive learning systems, XAI can help with individualized learning. By providing an explanation for the recommendations and assessments made by adaptive learning systems, XAI can help with individualized learning. We shall concentrate on a few particular applications in this section and go into detail about them.
3.1 Application of XAI in Medical Science
The field of artificial intelligence (AI) is rapidly growing on a global scale, particularly in healthcare, which is a hot topic for research [35]. There are numerous opportunities to utilize AI technology in the healthcare sector, where the well-being of individuals is at stake, due to its significant relevance and the vast amounts of digital medical data that have been collected [36]. AI has enabled us to perform tasks quickly that were previously unfeasible with traditional technologies.
The trustworthiness and openness of AI systems are becoming increasingly important, especially in areas like healthcare. As AI is used more in medical decision-making, people are worried about how reliable and understandable its results are. These worries highlight the need to evaluate AI models carefully to make sure their predictions are based on important and verifiable factors. In critical situations like medicine, proving that AI systems are credible is vital for their safe and effective use.
In the medical field, clinical decision support systems (CDSS) utilize AI technology to assist healthcare professionals with critical tasks such as diagnosis and treatment planning [37]. While these systems aim to support healthcare practitioners, misuse can have severe consequences in situations where lives are at risk. For example, false alarms, which are common in scenarios involving urgent patients, can lead to exhaustion among medical personnel.
The study [38] significantly contributes to medical skin lesion diagnostics in several ways. First, it modifies an existing explainable AI (XAI) technique to boost user confidence and trust in AI systems. This change involves developing an AI model that can distinguish between different types of skin lesions. The study uses synthetic examples and counter-examples to create explanations that highlight the key features influencing classification decisions. The research [38] trains a deep learning classifier with the ISIC 2019 dataset using the ResNet architecture. This allows professionals to use the explanations to reason effectively. Overall, the study’s main contributions lie in its refinement and evaluation of the XAI technique in a real-world medical setting, its analysis of the latent space, and its thorough user study to assess how effective the explanations are, particularly among experts in the field.
This research paper [39] discusses how to recognize brain tumors in MRI images using two effective algorithms: fuzzy C-means (FCM) and Artificial Neural Network (ANN). The authors aim to make the tumor segmentation process more understandable and improve accuracy in identifying tumors. Their main goal is to enhance tools that help doctors diagnose brain tumors more accurately.
This research offers two key benefits. First, it helps identify brain cancers in medical images more precisely, which is crucial for early diagnosis and treatment. Second, by incorporating XAI principles into the segmentation process, the researchers make their models’ decisions clearer and easier to understand for patients and medical experts. In summary, this increased clarity boosts the overall trust and acceptance of AI-driven systems in medical image analysis within clinical settings.
This study [40] discusses how AI and machine learning can help diagnose whole slide images (WSIs) in pathology. While AI can improve accuracy and efficiency, concerns exist about its reliability because it can be hard to understand. To address these issues, the article suggests using explainable AI methods, which help clarify how AI makes decisions. By adding XAI, pathology systems become more transparent and trustworthy, especially for critical tasks like diagnosing diseases. The study also introduces HistoMapr-Breast, a software tool that uses XAI to assist with breast core biopsies.
A recent study examines the importance of making sure AI systems in healthcare are accurate and strong, especially regarding how easy they are to understand and how well they can resist attacks [41]. As AI becomes more common in medical settings, it’s crucial to verify that the predictions these systems make rely on trustworthy features. To tackle this challenge, researchers have proposed various methods to improve model interpretability and explainability. The study shows that adversarial attacks can affect a model’s explainability, even when the model has strong training. Additionally, the authors introduce two types of attack classifiers: one that tells apart harmless and harmful inputs, and another that determines the nature of the attack.
This research paper [42] looks at explainable machine learning in cardiology. It discusses the challenges of understanding complex prediction models and how these models affect important healthcare decisions. The study explains the main ideas and methods of explainable machine learning, helping cardiologists understand the benefits and limitations of this approach. The goal is to improve decision-making in clinical settings by offering guidance on when to use easy-to-understand models versus complex ones. This can help improve patient outcomes while ensuring accountability and transparency in predictions.
Figure 9 shows a decision tree created from the predictions of a random forest model. This global tree diagram illustrates how the random forest works overall. By following a patient’s path through the tree, individual predictions can be examined. This type of explanation is beneficial because it clarifies both the general functioning of the model and the reasoning behind specific predictions. Decision trees are suitable for fields like cardiology because they use rule-based reasoning similar to clinical decision guidelines.
Figure 9: A model using random forests predicts heart disease by analyzing both local and global decision trees. The global diagram starts by examining whether a patient’s thallium stress test results are normal. If the test shows a problem, the model looks at the patient’s ST depression next. The local graphic shows the specific pathway a patient took in the model, explaining the reasons for their individual prediction. For example, a patient under 54.5 years old, with a maximum heart rate that is high and normal thallium stress test results, has a very low chance of having heart disease [42].
@@ 3.1 Application of XAI in Medical Science
Figure 10 shows the LIME explanations for our heart failure model’s two local predictions. The authors explain how these predictions serve as a clinical decision support tool in Epic, which is an electronic health record designed for doctors (Epic Systems Corporation, Verona, Wisconsin, USA). This kind of explanation helps to clarify the clinical factors that affect each prediction. Importantly, this type of explanation can be added to an EHR, which may improve the practical use of a complex model by making forecasts and clear explanations easy to integrate into clinical work.
@@
For medical AI to work reliably and be widely used, we need to do a lot of research and reach an agreement on important features like explainability, fairness, privacy, and reliability [33]. We must meet clear requirements and standards in any healthcare setting that uses AI, and we need to update these regularly. Additionally, we should establish laws that clarify who is responsible if something goes wrong with a medical AI whether that’s the designers, researchers, healthcare workers, or patients [43].
@@ 3.2 Explainability and Interpretability of Autonomous Systems
Explainability and interpretability are crucial concepts in the context of autonomous systems, referring to the ability to understand the decisions and behaviors of these systems. Explainability involves an autonomous system’s capacity to provide clear justifications for its actions and choices [44]. This clarity is essential for fostering acceptance and confidence in AI systems, especially in critical fields such as banking, healthcare, and autonomous vehicles.
While explainability and interpretability are closely related, interpretability focuses more on understanding the internal mechanisms and processes of the autonomous system [45]. An interpretable system offers users insight into the factors and criteria that influence its decision-making, enabling them to grasp how the system arrived at its conclusions.
@@
The research paper The research article [18] focuses on trust and dependability in autonomous systems. Autonomous systems have the potential for system operation, rapid information dissemination, massive data processing, working in hazardous environments, operating with greater resilience and tenacity than humans, and even astronomical examination [46], [47]. Following years of research and development, today’s automated technologies represent the peak of progress in computer recognition, responsive systems, user-friendly interface design, and sensing automation.
According to [44], the global market for automotive intelligent hardware, operations, and innovation is projected to grow significantly, increasing from $1.25billionin2017to$28.5 billion by 2025. Intel’s research on the expected benefits of autonomous vehicles indicates that implementing these technologies on public roads could reduce annual commute times by 250 million hours and save over 500,000 lives in the United States between 2035 and 2045 [44]. Modern cars utilize artificial intelligence for various functions, including intelligent cruise control, automatic driving and parking, and blind-spot detection (Figure 11).
@@
Authors [18] describe the challenges of autonomous systems, like, people sometimes tend to be overly excited about the potential of new ideas and ignore, or at least appear to be unaware of, the potential drawbacks of cutting-edge developments. Even in the early stages of robotics and autonomous system implementation, humanity preferred to put up with faulty goods and services, but they have gradually come to understand the importance of trustworthy and dependable autonomous systems. Numerous examples have demonstrated how operators’ use of automation is greatly impacted by trustworthiness.
@@
Figure 11: An automated vehicle that provides a clear and understandable rationale for its decisions at that moment serves as a prime example of explainable AI in automated driving [18].
@@
When AI has become prevalent in autonomous vehicle (AV) operations, user trust has been identified as a major issue that is essential to the success of these operations. XAI, which calls for the AI system to give the user explanations for every decision it makes, is a viable approach to fostering user trust for such integrated AI-based driving systems [48]. This work develops explainable Deep Learning (DL) models to improve trustworthiness in autonomous driving systems, driven by the need to improve user trust and the potential of innovative XAI technology in addressing such requirements. The main concept of this [48] research is to frame the decision-making process of autonomous vehicles (AVs) as an image-captioning task, generating textual descriptions of driving scenarios to serve as understandable explanations for humans. The proposed multi-modal deep learning architecture, shown in Figure 12, utilizes Transformers to model the relationship between images and language, generating meaningful descriptions and driving actions. Key contributions include improving the AV decision-making process for better explainability, developing a fully Transformer-based model, and outperforming baseline models. This results in enhanced user trust, valuable insights for AV developers, and improved interpretability through attention mechanisms and goal induction.
@@
This research [49] aims to investigate the integration of XAI into autonomous vehicular systems to improve transparency and human trust. It delves into the functioning of multiple inner vehicle modules, emphasizing the importance of understanding the vehicle’s decision-making processes for user credibility and reliability. The main contribution lies in introducing XAI to the domain of autonomous vehicles, showcasing its role in fostering trust, and highlighting advancements through comparative analysis. The output comprises the creation of visual explanatory techniques and an intrusion detection classifier, which show considerable advances over previous work in terms of transparency and safety in autonomous transportation systems.